OpenStudy (abbycross167):

Write the point-slope (y - k = m(x - h)) form of the equation from the information given. 1. Slope = ½; (h, k) = (1, -2) Can someone please show me how to do this?

OpenStudy (trantom):

so you know that m is the slope, so that part's easy then since h comes first before k in the ordered pair, you know that h is 1 and k is -2 so you plug those numbers into the equation. What do you get?

OpenStudy (trantom):

also (h,k) is a point which exists on the line given by the equation you get (just so you know what's going on)

OpenStudy (abbycross167):

-1? @trantom

OpenStudy (trantom):

you're writing the whole equation right?

OpenStudy (abbycross167):

ummm I guess so

OpenStudy (abbycross167):

I'm sorry i'm not good at math

OpenStudy (trantom):

its ok

OpenStudy (trantom):

just replace the letters m, k, and h with their values

OpenStudy (trantom):

m is the slope, and it says that slope=1/2 so you would replace m with 1/2

OpenStudy (trantom):

it says (h,k) = (1,-2) so you would replace h with 1 and replace k with -2

OpenStudy (abbycross167):

yes sir/ ma'am, would I subtract like this~ 1-2?

OpenStudy (trantom):

no those are separate values, the negative two and the positive 1 also I'm a guy xD

OpenStudy (abbycross167):

ok, haha so would it be -3? @trantom

OpenStudy (trantom):

this is just a point-slope equation, it doesn't have an "answer"

OpenStudy (trantom):

your answer would look like the y-k=m(x-h) but instead of k, m, and h you will have -2, 1/2 and 1

OpenStudy (trantom):

so for instance if you replace m, it would look like y-k=1/2(x-h)

OpenStudy (abbycross167):

ok umm if I show you an example from my lesson can you show me how to make it look like the example?

OpenStudy (trantom):

sure

OpenStudy (abbycross167):

Example 1: Find the equation of the line through the points (2, -1), (-4, 3) First, find the slope as we did in lesson 11. (y2 - y1) (x2 - x1) (3 - -1) (-4 - 2) (4) (-6) So our slope is -2/3 Next, pick one of the points (either one will do) to put into the equation y -k = m(x - h). Be sure to change the signs of the coordinates. y + 1 = -2/3(x - 2) or y -3 = -2/3(x + 4)

OpenStudy (abbycross167):

how do I make my answer look like that?

OpenStudy (trantom):

\[y-(-2)=\frac{ 1 }{ 2 }(x-1)\]

OpenStudy (trantom):

and since subtracting a negative is the same as adding, its really y+2

OpenStudy (trantom):

but see how I replaced each of those letters (m, h, and k) with their values?

OpenStudy (trantom):

remember m is the slope

OpenStudy (abbycross167):

yes sir, but this is kind of confusing :/

OpenStudy (trantom):

do you have math tomorrow at school?

OpenStudy (abbycross167):

yes sir

OpenStudy (trantom):

ok you might want to ask your math teacher or maybe a friend at school, I think an online chat can only go so far you know

OpenStudy (trantom):

I'll try one more thing

OpenStudy (trantom):



OpenStudy (abbycross167):

So the equation is y+2=1/2(x-1)?

OpenStudy (trantom):

yep!

OpenStudy (abbycross167):

wow. I'm really slow... thank you soooo much for helping me!!!

OpenStudy (trantom):

no problem :)